OpenStudy (anonymous):

iota(i)= root-1 and (i) square=1, how can we prove by (i).(i)= (i) square

OpenStudy (anonymous):

To be clear are you asking: Given \[\iota(i) = \sqrt{-1}\] and \[i^2 = 1\] prove \[i^2 = i * i\]

OpenStudy (anonymous):

the notation is confusing me a little...

OpenStudy (anonymous):

i = root(-1) so (i)*(i) = root(-1)*root(-1) = [root (-1)]^2 \[(\sqrt{-1})^{2} = -1\]

OpenStudy (anonymous):

I think in your question it shud be (i) square= -1 and NOT (i)square = 1

OpenStudy (anonymous):

yes... sorry sir...

OpenStudy (anonymous):

so (i)(i) = -1 = (i)^2

OpenStudy (anonymous):

no problem... here u just hv to prove LHS = RHS by the way kpelletij, where r u frm, I'm frm Delhi....

OpenStudy (anonymous):

sir i am frm delhi to!! living in gandhinagar..As my father is srvicing here.. hmm sir how root-1*root-1=1

OpenStudy (anonymous):

using the formula\[\sqrt{a}*\sqrt{b}=\sqrt{a*b}\] \[\sqrt{-1}*\sqrt{-1}= \sqrt{-1*-1}=\sqrt{1}=1\]

OpenStudy (anonymous):

@Harkirat sqrt(-1) * sqrt(-1) = i * i = i^2 = -1 Your formula is only valid for real non-negative a and b. Avoid the problem by converting to i at first opportunity.

OpenStudy (anonymous):

"how can we prove by (i).(i)= (i) square" You don't have to prove it, the definition of i is that i^2 = -1. Meaning it has square roots of plus and minus i, either of which squared = i^2.